UT 20131217 LUCI2 Nighttime Commissioning

Observer: JHeidt, APramskiy, MEdwards
Telescope Operator: GBechetti
LBTO Support: TSargent

Plan for the night

- MOS-acquisition
- - Further LS-tests, in particular Michelle practicing LS-acquisition

Summary

- MOS-acqusition needs to be optimized

- Did a good set of AGw-transform data

Preparations

Details

Times are in UT, files are luci2.20131216.iiii

00:20 Sunset: weather is moderate, sky covered with cirrus, T=8deg, hum =1\xDF%, low wind
had to skip flats

00:30 Tested a script to do offsets only (tel in track mode) for FF purposes

01:00 Telescope is collimated, good seeing, around 0.5"

MOS-acqusition

01:04 Preset to M225UD3H1, preset via Jaron's script failed, FW got stuck, recovered,
preset with own script, sky #1, scr #2, crap #3, throughslit #4, rotation matrix looks good, rms=0.05,0.06pix, applied offset, doesn' work #6,
Probably wrong rotangle has been applied (sign)
02:44 Redo, #8 sky, #9 src, #10 slit, compared with LS-acq. on same frames using just two stars,
results different (LS=-4.5,0.2,0.9, MOS=1.2,-4.7,2.4, for dex, dety, detrot),
changed the PA by 10degs, took image #11 and redid the aligment just using 2 stars
=> LS=1.2"-4",1deg, MOS=-4",25",11deg => Reults using LS look reasonable, MOS not (almost no change in rotation)

04:55 redo again, #13 sky, #14 src, #15 MOS, offsets: dx=-1.2", dy=-1.3", drot=2.14deg, #16-35 crap, #36 throughslit, no stars, #37 direct image, OK, systematic offset,
dx=-3.45, dy = 0.87", #38 wrong offset, changed sign in offset two times,
#39 => OK!!!! (subtract #39-#15 to see where stars are) => seems to work!!!
Some rotation still present most likely due to the center of rotation which is set to
x,y=1023,1023 but can reliably only be determined with a good AGw-transform

Some stats:
Refslit(#15),posslit(#15),posbluecirc(#15),posstar(#14),posalignstar(#17)
#31, 483/1284, 443/1291, 471/1280, 472/1288
#32, 480/937, 443/938, 481/925, 469/934
#34, 1730/1497, 1714/1516, 1736/1546, 1746,1506
#35, 1698/752, 1685/757, 1731/784, 1713/745

Refslit, diff posbluecirc-posalignstar, diff posslit-posalignstar #31, 11/-4, -29/3
#32, 11/3, -26/+4
#33, -16/-9, -32/10
#34, -15/7, -28/12

05:50 Preset to M225UD3H2, #40 sky, #41 src, #42 MOS, alignment using 6 stars
doesn't work, rms in detxy=17,4" !!! There seems something to be wrong with the mask.

AGw-transform data

06:40 Handover to DMiller for another AGW-transform set, seeing is excellent,ie worth to do it
Tried to do it simultaneously for LUCI1 and LUCI2, but gave up after 1 1/2 hours,
pointing on the right side was way off (ADSec was installed today).

08:20 Started to take AGw-transfrom data for LUCI2
Stone E1_15

STAR IMAGE

-----
BG 0049
46 0050
19 0051
10 0052
14 0053
30 0054
58 0055
40 0056
29 0057
11 0058
06 0059
09 0060
24 0061
31 0062
43 0063
05 0064
36 0065
39 0066
51 0067
35 0068
12 0069
22 0070
17 0071
25 0072
60 0073
62 0074
65 0075
53 FAIL (0076 was before collimation)
34 0077
45 0078
23 0079
07 0080
04 0081
08 0082
05 0083
36 0084
15 0085
26 0086

0, PA=0 0087 (on-axis with -x = 0 -y = 500)
0, PA=135 0088 (on-axis with -x = 0 -y = 500)
0, PA=180 0089 (on-axis with -x = 0 -y = 500)
0, PA=90 0090 (on-axis with -x = 0 -y = 500)
0, PA=90 0091 (on-axis with -x = 0 -y = 500)
0, PA=-90 0092 (on-axis with -x = 0 -y = 500)

MOS-acquisition

10:25 Preset to SDSS1110 for another MOS-alignment
#93 sky, #94 src, #95 MOS, field is misalinged, need to do the alignment first
Field alignment, x=-4300steps, image #95
11:10 Re-preset SDSS1110, #96 sky, #97 src, #98 MOS, offset dx=0.12, dy=-5.5, drot=3deg, looks reasonable, verification wo mos #99, off
Statistics:
slit, posslit, posalignedstar,diff
#31, 1010/1484, 1011/1462, 0/22
#32, 317/917, 300/909, 17/12
#33, 1010/1484, 1011/1462, 0/22

12:00 re-preset, failed, Geno reset TCS

12:18 re-preset, #100 sky, #101 src, #102 MOS,
doesn't either, learned from Ingo that new version is required

12:54 BS9142, throughput, done by Michelle, stopped without any notice after 4th spectrum
(z-band done completely), restarted, had a couple of GEIRS errors, stopped, #103-115

13:45 Moving telescope to zenith for flats, N3.75
H, 10*2.75s, #116-125, 10k
10*4.5s, #126-135, 18k
10*?, #136-145
K, 10*2.75s, #146-155, 8k
10*4.5s, #156-165, 15k
10*6s, #166-175, 22k
HeI, 10*2.75, #176-185, 14k
10*3.5s, #186-195, 25k
OH_1190, 10*2.75s, #196-205, 15k
10*3.5s, 10*3.5, #206-215, 22k
Br_gam, 10*2.7s, #216-225
10*7s, #226-235, 13k
10*7s, #236-245, 16k
10*10s, #246-255, 25K

14:20 Closing

-- JochenHeidt - 17 Dec2013
Topic revision: r7 - 17 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback